# Understanding the Dataset Practitioners Behind Large Language Model Development

### v1

@@ 7. Future work and conclusions
Through our retrospective analysis, we find that the dataset practitioner takes on a fluid role that is not well-defined in current literature on data workers. We hope that our contribution of defining this population and their use cases will enable the HCI community to better assess and support their needs.
In our interviews, we found that data quality is unanimously the top priority, but quality is subjective. Further research should explore what data quality means in different contexts, and how the same data can be high-quality or low-quality depending on the situation and perspective. Clarifying subjectivity across conceptual frameworks, evaluations, and workflows in this space remains a top priority.
Two primary data exploration patterns are identified: visually inspecting data in spreadsheets, which is not scalable, and writing custom analyses in Python notebooks, which is both high-effort and prone to confirmation bias. The community has not agreed upon alternative best-practices to explore this data, perhaps due to the newness of the space or the custom needs of the practitioners. There exist opportunities to better understand the landscape around both schools of thought: to formalize the lines of thinking in the shifting landscape, and to develop specific, flexible tooling solutions for custom analysis.
@@ Acknowledgements.
The authors wish to thank our colleagues at Google’s People + AI Research Team for helpful feedback and discussions, especially Michael Terry, Carrie Cai, and Michael Xieyang LiuReferences
@@
- Amershi et al. (2015) Saleema Amershi, Max Chickering, Steven M. Drucker, Bongshin Lee, Patrice Simard, and Jina Suh. 2015. ModelTracker: Redesigning Performance Analysis Tools for Machine Learning. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (CHI ’15). Association for Computing Machinery, New York, NY, USA, 337–346. https://doi.org/10.1145/2702123.2702509
@@
- Baldassarre et al. (2023) Maria Teresa Baldassarre, Danilo Caivano, Berenice Fernandez Nieto, Domenico Gigante, and Azzurra Ragone. 2023. The Social Impact of Generative AI: An Analysis on ChatGPT. In Proceedings of the 2023 ACM Conference on Information Technology for Social Good (GoodIT ’23). Association for Computing Machinery, New York, NY, USA, 363–373. https://doi.org/10.1145/3582515.3609555
- Bellamy et al. (2018) Rachel K. E. Bellamy, Kuntal Dey, Michael Hind, Samuel C. Hoffman, Stephanie Houde, Kalapriya Kannan, Pranay Lohia, Jacquelyn Martino, Sameep Mehta, Aleksandra Mojsilovic, Seema Nagar, Karthikeyan Natesan Ramamurthy, John T. Richards, Diptikalyan Saha, Prasanna Sattigeri, Moninder Singh, Kush R. Varshney, and Yunfeng Zhang. 2018. AI Fairness 360: An Extensible Toolkit for Detecting, Understanding, and Mitigating Unwanted Algorithmic Bias. CoRR abs/1810.01943 (2018). arXiv:1810.01943 http://arxiv.org/abs/1810.01943
@@
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 1877–1901. https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf
@@
- Chattopadhyay et al. (2020) Souti Chattopadhyay, Ishita Prasad, Austin Z. Henley, Anita Sarma, and Titus Barik. 2020. What’s Wrong with Computational Notebooks? Pain Points, Needs, and Design Opportunities. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3313831.3376729
@@
- Durrant-Whyte (2015) Hugh Durrant-Whyte. 2015. Data, Knowledge and Discovery: Machine Learning meets Natural Science. In Proceedings of the 21th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining (KDD ’15). Association for Computing Machinery, New York, NY, USA, 7. https://doi.org/10.1145/2783258.2785467
@@
- Han et al. (2020) Lei Han, Tianwa Chen, Gianluca Demartini, Marta Indulska, and Shazia Sadiq. 2020. On Understanding Data Worker Interaction Behaviors. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’20). Association for Computing Machinery, New York, NY, USA, 269–278. https://doi.org/10.1145/3397271.3401059
@@
- Harris et al. (2013) Harlan Harris, Sean Murphy, and Marck Vaisman. 2013. Analyzing the analyzers: An introspective survey of data scientists and their work. ” O’Reilly Media, Inc.”.
@@
- Kaur et al. (2020) Harmanpreet Kaur, Harsha Nori, Samuel Jenkins, Rich Caruana, Hanna Wallach, and Jennifer Wortman Vaughan. 2020. Interpreting Interpretability: Understanding Data Scientists’ Use of Interpretability Tools for Machine Learning. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3313831.3376219
- Kery et al. (2019) Mary Beth Kery, Bonnie E. John, Patrick O’Flaherty, Amber Horvath, and Brad A. Myers. 2019. Towards Effective Foraging by Data Scientists to Find Past Analysis Choices. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3290605.3300322
- Kery et al. (2018) Mary Beth Kery, Marissa Radensky, Mahima Arya, Bonnie E. John, and Brad A. Myers. 2018. The Story in the Notebook: Exploratory Data Science using a Literate Programming Tool. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–11. https://doi.org/10.1145/3173574.3173748
- Kim et al. (2018) Been Kim, Martin Wattenberg, Justin Gilmer, Carrie Cai, James Wexler, Fernanda Viegas, and Rory sayres. 2018. Interpretability Beyond Feature Attribution: Quantitative Testing with Concept Activation Vectors (TCAV). In Proceedings of the 35th International Conference on Machine Learning (Proceedings of Machine Learning Research), Jennifer Dy and Andreas Krause (Eds.), Vol. 80. PMLR, 2668–2677. https://proceedings.mlr.press/v80/kim18d.html
- Li et al. (2023) Catherine Li, Talie Massachi, Jordan Eschler, and Jeff Huang. 2023. Understanding the Needs of Enterprise Users in Collaborative Python Notebooks: This paper examines enterprise user needs in collaborative Python notebooks through a dyadic interview study. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (CHI EA ’23). Association for Computing Machinery, New York, NY, USA, Article 402, 7 pages. https://doi.org/10.1145/3544549.3573843
@@
- Mehrabi et al. (2021) Ninareh Mehrabi, Fred Morstatter, Nripsuta Saxena, Kristina Lerman, and Aram Galstyan. 2021. A Survey on Bias and Fairness in Machine Learning. ACM Comput. Surv. 54, 6, Article 115 (jul 2021), 35 pages. https://doi.org/10.1145/3457607
@@
- Muller et al. (2019) Michael Muller, Ingrid Lange, Dakuo Wang, David Piorkowski, Jason Tsay, Q. Vera Liao, Casey Dugan, and Thomas Erickson. 2019. How Data Science Workers Work with Data: Discovery, Capture, Curation, Design, Creation. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3290605.3300356
- OpenAI (2023) R OpenAI. 2023. Gpt-4 technical report. arxiv 2303.08774. View in Article 2 (2023), 13.
@@
- Sambasivan et al. (2021) Nithya Sambasivan, Shivani Kapania, Hannah Highfill, Diana Akrong, Praveen Paritosh, and Lora M Aroyo. 2021. “Everyone wants to do the model work, not the data work”: Data Cascades in High-Stakes AI. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 39, 15 pages. https://doi.org/10.1145/3411764.3445518
- Tabard et al. (2008) Aurélien Tabard, Wendy E. Mackay, and Evelyn Eastmond. 2008. From individual to collaborative: the evolution of prism, a hybrid laboratory notebook. In Proceedings of the 2008 ACM Conference on Computer Supported Cooperative Work (CSCW ’08). Association for Computing Machinery, New York, NY, USA, 569–578. https://doi.org/10.1145/1460563.1460653
@@
- Wang et al. (2019) Dakuo Wang, Justin D. Weisz, Michael Muller, Parikshit Ram, Werner Geyer, Casey Dugan, Yla Tausczik, Horst Samulowitz, and Alexander Gray. 2019. Human-AI Collaboration in Data Science: Exploring Data Scientists’ Perceptions of Automated AI. Proc. ACM Hum.-Comput. Interact. 3, CSCW, Article 211 (nov 2019), 24 pages. https://doi.org/10.1145/3359313
@@
![[Uncaptioned image]](extracted/5423847/results_table.png)

### v2

@@ 6. Conclusions and Future Work
Through our retrospective analysis, we found that the dataset practitioner takes on a fluid role that is not well-defined in current literature on data workers. We hope that our contribution of defining this population and their use cases will enable the HCI community to better assess and support their needs.
In our interviews, we found that data quality is unanimously the top priority, but quality is subjective. Further research should explore what data quality means in different contexts, and how the same data can be high-quality or low-quality depending on the situation and perspective. Clarifying subjectivity across conceptual frameworks, evaluations, and workflows in this domain remains a top priority, potentially achieved through standardizing metrics (e.g. toxicity, distributions of relevant safety features, data diversity) and evaluation criteria.
Two primary data exploration patterns emerge: visually inspecting data in spreadsheets, which lacks scalability, and crafting custom analyses in Python notebooks, which is high-effort. Both practices are susceptible to confirmation bias. However, the community has yet to reach a consensus on alternative best-practices to for data exploration, possibly due to the nascent nature of the field or the custom needs of the practitioners. There are opportunities to determine the specific areas where prioritizing either flexibility or specificity is most beneficial; these opportunities can be addressed by formalizing evaluation frameworks in the evolving landscape, and developing flexible tooling for custom analysis.
@@ Acknowledgements.
The authors wish to thank our study participants and Google’s People + AI Research Team (PAIR), especially James Wexler and Michael Terry.References
@@
- Amershi et al. (2015) Saleema Amershi, Max Chickering, Steven M. Drucker, Bongshin Lee, Patrice Simard, and Jina Suh. 2015. ModelTracker: Redesigning Performance Analysis Tools for Machine Learning. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (CHI ’15). ACM, 337–346. https://doi.org/10.1145/2702123.2702509
@@
- Ashtari et al. (2023) Narges Ashtari, Ryan Mullins, Crystal Qian, James Wexler, Ian Tenney, and Mahima Pushkarna. 2023. From Discovery to Adoption: Understanding the ML Practitioners’ Interpretability Journey. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (DIS ’23). Association for Computing Machinery, New York, NY, USA, 2304–2325. https://doi.org/10.1145/3563657.3596046
- Baldassarre et al. (2023) Maria Teresa Baldassarre, Danilo Caivano, Berenice Fernandez Nieto, Domenico Gigante, and Azzurra Ragone. 2023. The Social Impact of Generative AI: An Analysis on ChatGPT. In Proceedings of the 2023 ACM Conference on Information Technology for Social Good (GoodIT ’23). ACM, 363–373. https://doi.org/10.1145/3582515.3609555
- Bellamy et al. (2018) Rachel K. E. Bellamy, Kuntal Dey, Michael Hind, Samuel C. Hoffman, Stephanie Houde, Kalapriya Kannan, Pranay Lohia, Jacquelyn Martino, Sameep Mehta, Aleksandra Mojsilovic, Seema Nagar, Karthikeyan Natesan Ramamurthy, John T. Richards, Diptikalyan Saha, Prasanna Sattigeri, Moninder Singh, Kush R. Varshney, and Yunfeng Zhang. 2018. AI Fairness 360: An Extensible Toolkit for Detecting, Understanding, and Mitigating Unwanted Algorithmic Bias. arXiv preprint arXiv:1810.01943 (2018). http://arxiv.org/abs/1810.01943
- Birch et al. (2018) David Birch, David Lyford-Smith, and Yike Guo. 2018. The Future of Spreadsheets in the Big Data Era. arXiv:cs.CY/1801.10231
@@
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. In Advances in Neural Information Processing Systems, Vol. 33. 1877–1901. https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf
@@
- Chattopadhyay et al. (2020) Souti Chattopadhyay, Ishita Prasad, Austin Z. Henley, Anita Sarma, and Titus Barik. 2020. What’s Wrong with Computational Notebooks? Pain Points, Needs, and Design Opportunities. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). ACM, 1–12. https://doi.org/10.1145/3313831.3376729
@@
- Durrant-Whyte (2015) Hugh Durrant-Whyte. 2015. Data, Knowledge and Discovery: Machine Learning meets Natural Science. In Proceedings of the 21th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining (KDD ’15). ACM, 7. https://doi.org/10.1145/2783258.2785467
@@
- Harris et al. (2013) Harlan Harris, Sean Murphy, and Marck Vaisman. 2013. Analyzing the analyzers: An introspective survey of data scientists and their work. O’Reilly Media, Inc.
@@
- Hohman et al. (2018) Fred Hohman, Minsuk Kahng, Robert Pienta, and Duen Horng Chau. 2018. Visual analytics in deep learning: An interrogative survey for the next frontiers. IEEE Transactions on Visualization and Computer Graphics 25, 8 (2018), 2674–2693. https://doi.org/10.1109/TVCG.2018.2843369
- Kahng et al. (2024) Minsuk Kahng, Ian Tenney, Mahima Pushkarna, Michael Xieyang Liu, James Wexler, Emily Reif, Krystal Kallarackal, Minsuk Chang, Michael Terry, and Lucas Dixon. 2024. LLM Comparator: Visual Analytics for Side-by-Side Evaluation of Large Language Models. Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems.
@@
- Kaur et al. (2020) Harmanpreet Kaur, Harsha Nori, Samuel Jenkins, Rich Caruana, Hanna Wallach, and Jennifer Wortman Vaughan. 2020. Interpreting Interpretability: Understanding Data Scientists’ Use of Interpretability Tools for Machine Learning. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (CHI ’20). ACM, 1–14. https://doi.org/10.1145/3313831.3376219
- Kery et al. (2019) Mary Beth Kery, Bonnie E. John, Patrick O’Flaherty, Amber Horvath, and Brad A. Myers. 2019. Towards Effective Foraging by Data Scientists to Find Past Analysis Choices. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). ACM, 1–13. https://doi.org/10.1145/3290605.3300322
- Kery et al. (2018) Mary Beth Kery, Marissa Radensky, Mahima Arya, Bonnie E. John, and Brad A. Myers. 2018. The Story in the Notebook: Exploratory Data Science using a Literate Programming Tool. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). ACM, 1–11. https://doi.org/10.1145/3173574.3173748
- Kim et al. (2018) Been Kim, Martin Wattenberg, Justin Gilmer, Carrie Cai, James Wexler, Fernanda Viegas, and Rory sayres. 2018. Interpretability Beyond Feature Attribution: Quantitative Testing with Concept Activation Vectors (TCAV). In Proceedings of the 35th International Conference on Machine Learning, Vol. 80. PMLR, 2668–2677. https://proceedings.mlr.press/v80/kim18d.html
- La Rosa et al. (2023) Biagio La Rosa, Graziano Blasilli, Romain Bourqui, David Auber, Giuseppe Santucci, Roberto Capobianco, Enrico Bertini, Romain Giot, and Marco Angelini. 2023. State of the art of visual analytics for explainable deep learning. In Computer Graphics Forum, Vol. 42. Wiley Online Library, 319–355.
- Li et al. (2023) Catherine Li, Talie Massachi, Jordan Eschler, and Jeff Huang. 2023. Understanding the Needs of Enterprise Users in Collaborative Python Notebooks: This paper examines enterprise user needs in collaborative Python notebooks through a dyadic interview study. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (CHI EA ’23). ACM, Article 402, 7 pages. https://doi.org/10.1145/3544549.3573843
@@
- Mehrabi et al. (2021) Ninareh Mehrabi, Fred Morstatter, Nripsuta Saxena, Kristina Lerman, and Aram Galstyan. 2021. A Survey on Bias and Fairness in Machine Learning. Comput. Surveys 54, 6, Article 115 (2021), 35 pages. https://doi.org/10.1145/3457607
@@
- Muller et al. (2019) Michael Muller, Ingrid Lange, Dakuo Wang, David Piorkowski, Jason Tsay, Q. Vera Liao, Casey Dugan, and Thomas Erickson. 2019. How Data Science Workers Work with Data: Discovery, Capture, Curation, Design, Creation. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). ACM, 1–15. https://doi.org/10.1145/3290605.3300356
- Nahar et al. (2022) Nadia Nahar, Shurui Zhou, Grace Lewis, and Christian Kästner. 2022. Collaboration challenges in building ML-enabled systems: communication, documentation, engineering, and process. In Proceedings of the 44th International Conference on Software Engineering (ICSE ’22). ACM, 413–425. https://doi.org/10.1145/3510003.3510209
- OpenAI (2023) OpenAI. 2023. GPT-4 technical report. arXiv preprint arXiv:2303.08774 (2023).
@@
- Reif et al. (2024) Emily Reif, Crystal Qian, James Wexler, and Minsuk Kahng. 2024. Automatic Histograms: Leveraging Language Models for Text Dataset Exploration. Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems.
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. ”Why Should I Trust You?”: Explaining the Predictions of Any Classifier. In Proceedings of the 22nd ACM SIGKDD International Conference on Knowledge Discovery and Data Mining (KDD ’16). ACM, 1135–1144. https://doi.org/10.1145/2939672.2939778
- Sambasivan et al. (2021) Nithya Sambasivan, Shivani Kapania, Hannah Highfill, Diana Akrong, Praveen Paritosh, and Lora M Aroyo. 2021. “Everyone wants to do the model work, not the data work”: Data Cascades in High-Stakes AI. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). ACM, Article 39, 15 pages. https://doi.org/10.1145/3411764.3445518
- Tabard et al. (2008) Aurélien Tabard, Wendy E. Mackay, and Evelyn Eastmond. 2008. From individual to collaborative: the evolution of prism, a hybrid laboratory notebook. In Proceedings of the 2008 ACM Conference on Computer Supported Cooperative Work (CSCW ’08). ACM, 569–578. https://doi.org/10.1145/1460563.1460653
- Team et al. (2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Yonghui Wu, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, et al. 2023. Gemini: a family of highly capable multimodal models. arXiv preprint arXiv:2312.11805 (2023).
@@
- Wang et al. (2019) Dakuo Wang, Justin D. Weisz, Michael Muller, Parikshit Ram, Werner Geyer, Casey Dugan, Yla Tausczik, Horst Samulowitz, and Alexander Gray. 2019. Human-AI Collaboration in Data Science: Exploring Data Scientists’ Perceptions of Automated AI. Proceedings of the ACM on Human-Computer Interaction 3, CSCW, Article 211 (2019), 24 pages. https://doi.org/10.1145/3359313
@@
- Zhang et al. (2020) Amy X. Zhang, Michael Muller, and Dakuo Wang. 2020. How do Data Science Workers Collaborate? Roles, Workflows, and Tools. Proceedings of the ACM on Human-Computer Interaction 4, CSCW1 (2020), 1–23.